OpenStudy (anonymous):

Use the following to determine the distance A (13,7) B (-9,4). 1) 5 2) √493 3) √137 4) √112

OpenStudy (anonymous):

distance equals \[\sqrt{(4-7)^{2}+(-9-13)^{2}}=\sqrt{493}\]

OpenStudy (anonymous):

Use the distance formula: \[d = \sqrt{(x_{1} - x_{2})^{2} + (y_{1} - y_{2})^{2}}\]

OpenStudy (campbell_st):

draw a picture and then find some lengths  a right triangle is formed... used pythagoras' theorem to find d d^2 = 22^2 + 3^2 \[d = \sqrt{493}\]